I do not have enough savings now to buy pair of shoes.If I increase my savings by 15%, I would still need another 9$.If I increase my savings by 35% , I would have 17$ more than I need.How much money do I have in my savings now.
Answer by ikleyn(52835)

.
Let x = How much money you have in your savings now.

Let y = the cost of the pair of shes.


Then from the condition you have these two equations


    x + 0.15x = y -  9     dollars     (1)

    x + 0.35x = y + 17     dollars     (2)


Subtract equation (1) from equation (2) to eliminate x. You will get

       0.35x - 0.15x = 17 - (-9)

       0.2x          = 26

       x             = 26/0.2 = 130 dollars.


ANSWER.  Your savings now is/(are) 130 dollars.

Solved.